Documenta Mathematica, Journal der Deutschen Mathematiker-Vereinigung, publishes research manuscripts out of all mathematical fields and is refereed in the traditional manner.

Documenta Mathematica is published on the World Wide Web under the address:

          http://www.mathematik.uni-bielefeld.de/documenta

Manuscripts should be submitted as TEX files by e-mail to one of the editors. Hints for manuscript preparation can be found under the above WWW-address.

Managing Editors:

Alfred K. Louis, Saarbrücken  louis@num.uni-sb.de
Ulf Rehmann (techn.), Bielefeld  rehmann@mathematik.uni-bielefeld.de
Peter Schneider, Münster  pschnei@math.uni-muenster.de

Editors:

Don Blasius, Los Angeles  blasius@math.ucla.edu
Joachim Cuntz, Heidelberg  cuntz@math.uni-muenster.de
Bernold Fiedler, Berlin (FU)  fiedler@math.fu-berlin.de
Friedrich Götze, Bielefeld  goetze@mathematik.uni-bielefeld.de
Wolfgang Hackbusch, Kiel  wh@informatik.uni-kiel.de
Ursula Hamenstädt, Bonn  ursula@rhein.iam.uni-bonn.de
Max Karoubi, Paris  karoubi@math.jussieu.fr
Rainer Kreß, Göttingen  kress@math.uni-goettingen.de
Stephen Lichtenbaum, Providence  Stephen Lichtenbaum@brown.edu
Alexander S. Merkurjev, Los Angeles merkurev@math.ucla.edu
Anil Nerode, Ithaca  anil@math.cornell.edu
Thomas Peternell, Bayreuth  peternel@btm8x1.mat.uni-bayreuth.de
Wolfgang Soergel, Freiburg  soergel@sun2.mathematik.uni-freiburg.de
Günter M. Ziegler, Berlin (TU)  ziegler@math.tu-berlin.de

Editors of the Extra Volume ICM 1998: Ceremonies, Laudationes, Prize Winners, Photos:

Gerd Fischer, Düsseldorf  gerdfischer@cs.uni-duesseldorf.de
Hannes Stoppel (Editorial Assistant) stoppel@cs.uni-duesseldorf.de

Plenary and Invited Lectures:

Ulf Rehmann, Bielefeld  rehmann@mathematik.uni-bielefeld.de

This volume was produced electronically, using various dialects of TEX, at the Fakultät für Mathematik, University of Bielefeld, Germany. \begin{matrix}c \\ ◯ \end{matrix} 1998 for Layout: Ulf Rehmann, Technical Managing Editor, Fakultät für Mathematik, Universität Bielefeld, Postfach 100131, D-33501 Bielefeld, Germany.

ISSN 1431-0635 Print ISSN 1431-0643 Internet

Print: Publishing On Demand by Geronimo GmbH, D-83026 Rosenheim, Germany

                       http://www.geronimo.de